A tuned radio frequency has the ff. components: C=12 pF, L=9.6 mH, RL=320 Ω.        The TRF is tuned to 520 kHz. Determine the bandwidth, quality factor and dissipation factor.
Answer #1

Given : C=12 pF, L=9.6 mH, R=320 Ω.and fr =520 KHz,

The TRF receiver is a RARALLEL RLC CIRCUIT resonating at fr Hz.

1. Quality factor Q= R/2\pifr L

Q= 320 / (2x3.14x520x10^3 x9.6x10^-3) = 0.0102.

2. Bandwidth= fr / Q = 520x10^3 / 0.012 = 50.969200 GHz.

3. Dissipation factor is inverse of quality factor.

\delta =1/Q = 1 / 50.969200x 10^6 = 74.81 .